A stone used in the sport of curling has a mass of 18.0 kg and is initially at rest, sitting on a

Flat ice surface. It is pushed with a constant force of magnitude 24.0N over a distance of

4.00 m. (You may assume that any friction between the stone and the ice is negligible.)

(a) (i) What is the acceleration of the stone and what is the work done in accelerating it?

(ii) By considering the conservation of energy of the stone, calculate the final speed of the

stone. Show that you have checked that the value for the magnitude of speed and its unit are

Sensible.

(b) How much power is supplied by the person pushing the stone at the beginning of the

Push, and how much is supplied at the end of the push?

(c) Calculate the duration of the push from the acceleration and final speed. Compare this

with the time calculated from the average power supplied during the push and the amount of

work done
Answer #1

224x4-96 - a = a) sa = 4y => 1-33 ntje W=24x47960 (81) AkE- Wordone. ¿(v2-42) = 24x4 am kg maxim s ļ (v2.0)=24x4 kg m/ ² v=12